# TaCo: Targeted Concept Removal in Output Embeddings for NLP via Information Theory and Explainability

### v1

@@ 4 Methodology
Now that we’ve established the constraints on how to implement our method, we define the tools used for each part of the method: (i) For the decomposition that uncovers our concepts, we use the SVD decomposition. (ii) For the gender importance calculation, we train a classifier for this task and then use Sobol importance. (iii) To remove concepts, we simply remove the columns corresponding to the removed concept.
@@ 4.1 Generating concepts used for occupation classification
Firstly, we look into latent space before the last layer of the model, where the decisions are linear. We have an output embedding matrix (also called CLS embedding) of the transformer model in this space. The aim here is to employ decomposition techniques to extract both a concept coefficients matrix and a concept base matrix from latent vectors.
To illustrate this part, we discover concepts without supervision by factorizing the final embedding matrix through a Singular Value Decomposition. We choose this technique, but it is possible to do other decompositions, like PCA Wold et al. (1987), RCA Bingham and Mannila (2001) or ICA Comon (1994).
Singular Value Decomposition (SVD)
The embedding matrix $\bm{A}\in\mathbb{R}^{n\times d}$ can be decomposed using an SVD as:
$\bm{A}=U_{0}\Sigma_{0}V_{0}^{\intercal}=U_{0}W_{0}\,,$
where $U_{0}\in\mathbb{R}^{n\times d}$ contains the left singular vectors, $V_{0}\in\mathbb{R}^{d\times d}$ contains the right singular vectors, and $\Sigma_{0}\in\mathbb{R}^{d\times d}$ is a diagonal matrix containing the singular values. Note that $U_{0}$ and $V_{0}$ are orthonormal matrices, which means that $U_{0}^{\intercal}U_{0}=Id$ and $V_{0}^{\intercal}V_{0}=Id$.
An SVD can be used to explain the main sources of variability of the decomposed matrix, here $\bm{A}$. The matrix $W_{0}^{-1}=V_{0}\Sigma_{0}^{-1}$ is indeed considered as a linear projector of $\bm{A}$, and $U_{0}$ is the projection of $\bm{A}$ on a basis defined by $V_{0}$. The amount of variability, or more generally, of information that has been captured in $\bm{A}$ by a column of $U_{0}$ is directly represented by the corresponding singular value in $\Sigma$. Using only the $r$ largest singular values of $\Sigma_{0}$, with $r<<d$, therefore allows capturing as much variability in $\bm{A}$ as possible with a reduced amount of dimensions. We then decompose $\bm{A}$ as:
@@
where $\bm{W}=\Sigma V^{\intercal}$ is a compressed version of $W_{0}$, in the sense that $\Sigma\in\mathbb{R}^{r,r}$ is the diagonal matrix containing the $r$ largest singular values of $\Sigma$ and $V\in\mathbb{R}^{d,r}$ contains the $r$ corresponding right singular vectors. The matrix $\bm{U}\in\mathbb{R}^{n,r}$ consequently contains the $r$ right singular vectors which optimally represent the information contained in $\bm{A}$ according to the SVD decomposition. In this paper, we refer to these vectors as the extracted concepts of $\bm{A}$. Python implementation details are in Appendix D.
@@ 4.2 Estimating concept importance with Sobol Indices
We then evaluate the concepts’ importance by applying a feature importance technique on $\bm{G}$ using $c_{g}$, and on $\bm{Y}$ using $c$. More specifically, Sobol indices Sobol (1993) are used to estimate the concepts’ importance, as in Jourdan et al. (2023). A key novelty of the methodology developed in our paper is to simultaneously measure each concept importance for the outputs $\bm{G}$ and $\bm{Y}$, and not only $\bm{Y}$ as proposed in Jourdan et al. (2023).
Figure 2: Co-importance plot for $r=20$ dimensions with respect to Occupation and Gender labels, computed with the Sobol method. The colour is based on the angle $a=90-\frac{2}{\pi}\arctan{(\frac{y}{x})}$. The angles with extreme values correspond to concepts with high importance for gender, but comparatively low importance for occupation. Their removal yields a favourable tradeoff in accuracy/fairness.
To estimate the importance of a concept $\bm{U}_{i}$ for gender (resp. for the label), we measure the fluctuations of the model output $c_{g}(\bm{U}\bm{W})$ (resp. $c(\bm{U}\bm{W})$) in response to perturbations of concept coefficient $\bm{U}_{i}$
We then propagate this perturbed activation to the model output $\bm{G}=c_{g}(\tilde{\bm{A}})$ (resp. $\bm{Y}=c(\tilde{\bm{A}})$). We can capture the importance that a concept might have as a main effect – along with its interactions with other concepts – on the model’s output by calculating the expected variance that would remain if all the concepts except the i were to be fixed. This yields the general definition of the total Sobol indices. An important concept will have a large variance in the model’s output, while an unused concept will barely change it.
We explain the total Sobol indices for each concept in details in Appendix E, along with our perturbation strategy.
@@ 4.3 Neutralizing the embedding with importance-based concept removal
In this final section, we propose to eliminate $k$ concepts that exhibit significant gender dependence within the final embedding matrix, with the objective of improving its gender-neutrality. This provides us with an improved embedding which can serve as a base to learn a fairer classifier to predict downstream tasks, and thus, a more gender-neutral model overall.
This concept-removal procedure consists of choosing the $k$ concepts that maximize the ratio between the importance for gender prediction and for the task. This graphically corresponds to searching for the concepts whose angle with respect to the line of equal importance for gender and task prediction are the highest (see Fig. 2). This strategy allows us to work on a trade-off between the model’s performance and its fairness.
In practice, we sort the concepts $\bm{U}_{1}$,$\bm{U}_{2}$,….$\bm{U}_{r}$ according to the second strategy in descending order and we delete the $k$ first concepts. This essentially leaves us with a concept base $\tilde{\bm{U}}=(\bm{U}_{k+1},...,\bm{U}_{r})$. This same transformation can be applied to $\bm{W}$’s lines, yielding $\tilde{\bm{W}}$. From the matrix obtained from the matrix product $\tilde{\bm{U}}\tilde{\bm{W}}$, we retrain a classifier $\tilde{c}$ to predict $\bm{Y}$ (refer to Appendix C.2 for more details).
Remark:
The more concepts we remove, the less we will be able to predict the gender, so the fairer the model will be. However, this also entails a decrease on the amount of information that is encoded in the embedding and that can be used for downstream tasks. The parameter $k$ must therefore be chosen to comply with our specific application’s constraints.
Figure 3: Accuracy drop when $k$ concepts are removed for an SVD with $r=20$ dimensions. In Figure 2(a), (Top) the occupation accuracy, (Bottom) the gender accuracy. Figure 2(b) shows occupation accuracy as a function of gender accuracy drop. The gender accuracy drop is reported relatively to the Baseline1, the classifier trained on Bios with all the concepts.
@@ 5.1 Dataset and models
To illustrate our method, we use the Bios dataset De-Arteaga et al. (2019), which contains about 440K biographies (textual data) with labels for the genders (M or F) of the authors, and their occupations (28 classes). From the Bios dataset, we also create a second dataset Bios-neutral by removing all the explicit gender indicators. Bios-neutral will be used as a baseline to which we can compare our method. We leave the details of how this dataset preprocessing was performed in Appendix B.
For both datasets, we train a RoBERTa (Liu et al., 2019a) model to predict the occupation task. We leave the details on how these models were trained in Appendix C.
5.2 Results analysis - Quantitative part
Figure 4: Understanding concepts extracted from the model trained on Bios. Concept 1 prominently features explicit gender indicators like ’she’ and ’his’, crucial for gender identification but irrelevant for occupation. Concept 8 integrates information about gender-unbalanced professions, aiding in predicting both, an individual’s gender and occupation. Concept 19 focuses on information related to the professor occupation, a gender-balanced class and most represented in the dataset.
Initially, we will examine the progression of our model’s performance for each dimension removed (one by one) utilizing our methodology.
Following the precepts of information theory, we will assess the effectiveness of our bias mitigation strategy by measuring the model’s capacity to predict the gender and the task as we progressively remove concepts.
We showcase in Fig. 3 the outcome of removing the $k$ most important concepts for $k\in\llbracket 0,19\rrbracket$ (we have $r=20$ dimensions) in terms of different metrics: accuracy on gender prediction and on predicting occupation (downstream task). We compare our technique to two baselines: Baseline1 – RoBERTa model trained on Bios dataset (i.e. when $k=0$) –, and Baseline2 – RoBERTa model trained on Bios-neutral dataset.
From Figure 3, simply by removing one concept, we achieve significant gains in the gender neutrality of our model, as it predicts gender with much lower accuracy (82.3% compared to the initial 96%) while maintaining its performance in occupation prediction (accuracy decreasing from 86.4% to 86.3%). Starting from the removal of the fourth dimension, the accuracy of gender prediction falls below the threshold of Baseline2, with 79.4% accuracy compared to 79.8% for Baseline2, while maintaining an accuracy of 86.3% for occupation (compared to 86% for Baseline2). Moreover, our method, in addition to being explainable, time-efficient, and cost-effective, exhibits greater effectiveness.
After removing these first four concepts, we enter a regime where removing concepts continues to gain us fairness (as the accuracy of the gender task continues to fall), while the accuracy of the occupation remains unchanged. This regime is therefore the one with the best trade-off between accuracy and fairness (shown in green here). From concept 12 onwards, the accuracy of occupation begins to decline, as we look more and more at concepts with a high importance for occupation (according to Sobol), and we enter a new regime where we gain more and more in fairness, but where the accuracy of our task collapses. It may be worthwhile deleting a fairly large number of concepts to get into this zone, but only if we are willing to pay the price in terms of prediction quality.
It is important to note that in this specific application, the optimal fairness-accuracy trade-off corresponds to accuracy on gender prediction of 62% as that’s the accuracy one gets when learning a model to predict the gender with the occupation labels as only information.
5.3 Concepts removed analysis - Explainability part
In this section, we interpret the dimensions of the Singular Value Decomposition (SVD) as valuable concepts for predicting our two tasks: gender and occupation. For this, we use the interpretability part of the COCKATIEL Jourdan et al. (2023) method on concepts/dimensions previously found with the SVD and ranked with the Sobol method by importance for gender and occupation.
Concept interpretability
COCKATIEL modifies the Occlusion method Zeiler and Fergus (2014), which operates by masking each word and subsequently observing the resultant impact on the model’s output. In this instance, to infer the significance of each word pertaining to a specific concept, words within a sentence are obscured, and the influence of the modified sentence (devoid of the words) on the concept is measured. This procedure can be executed at either the word or clause level, meaning it can obscure individual words or entire clauses, yielding explanations of varying granularity contingent upon the application.
Concepts interpretation
For the dimensions of the SVD to be meaningful for explanation, they must first be meaningful for the model’s prediction, whether it be the model predicting gender or the one predicting occupation. We rank the concepts w.r.t their removal order, and we focus on three important ones:
Concept 1, with a gender importance of 32%, and an occupation importance of 5%; Concept 8, with a gender importance of 16%, and an occupation importance of 12%; Concept 19, with a gender importance of 12%, and an occupation importance of 33%.
These three concepts are crucial for the prediction of at least one of our two tasks. Moreover, they are representative of our problem statement as Concept 1 is more important for gender, Concept 8 is as important to gender as occupation, and Concept 19 is more important for occupation. We illustrate several examples of explanations for these three concepts in Figure 4 to enable qualitative analysis.
According to the importance per dimension calculated in part 2 of our method, concept 1 is almost exclusively predictive of gender. The most significant words for the concept of this dimension are explicit indicators of traditional gender.
Concept 8 is capable of predicting both gender and occupation. Within its concept, we find examples of biographies predominantly related to unbalanced professions such as dietitian (92.8% women), chiropractor (26.8% women), software engineer (16.2% women), and model (83% women). It is logical that the concept allowing the prediction of these professions also enables the prediction of gender, as it is strongly correlated in these particular professions.
Concept 19, which strongly predicts occupation, focuses on biographies predicting professors (constituting more than 29.7% of all individuals). This class is gender-balanced (with 45.1% women), providing little information on gender.
@@ 6 Conclusion
In this paper, we introduced a method for neutralizing the presence of bias in transformer model’s embeddings in an interpretable and cost-efficient way through concept-based explainability. In this study, we present a method applied to gender-related information, yet inherently adaptable to other sensitive variables. Indeed, by removing concepts that are mostly influential for gender prediction, we obtain embeddings that rely considerably less on gender information for generating predictions, and at the same time, remain highly accurate at the task. Furthermore, by doing so via explainable AI, we perform interventions that can be easily interpreted by humans, and thus gain trust in the bias mitigation process and insights into what produced the bias in the first place. Finally, thanks to its low computation cost, it can be easily applied to pre-trained, fine-tuned and even embeddings accessed through inference APIs, thus allowing our method to be applied to any model in a post-hoc manner.
@@
Herein, we propose a novel approach that serves as a framework that can be applied to various concept extraction techniques and feature importance computation methods. This means that the SVD could be exchanged for other dimensionality reduction techniques (and thus, change the properties of the concept base and coefficients), or Shapley values could be preferred over Sobol indices for the relaxation on the independence assumption, for istance. In addition, our application was gender bias mitigation, but it could be applied to any other type of bias as long as the sensitive variable is available or can be estimated.

### v4

@@ 4 Methodology
Now that we’ve established constraints on how to implement our method, we define the tools used for each part of the method. (i) For the decomposition that uncovers our concepts, we use different decompositions: SVD, PCA, and ICA. (ii) For the sensitive variable importance calculation, we train a classifier for this task and then use Sobol importance. (iii) To remove concepts, we simply remove the columns of the matrix that correspond to the target concept.
@@ 4.1 Generating concepts used for occupation classification
Firstly, we examine the latent space before the last layer of the model. We have an final latent embedding $\bm{A}\in\mathbb{R}^{n\times d}$ of the transformer model in this space. The aim here is to employ matrix decomposition techniques to extract both a concept coefficients matrix $\bm{U}$ and a concept base matrix $\bm{W}$ for the latent vectors.
To illustrate this part, we discover concepts without supervision by factorizing the final embedding matrix using different decompositions: singular value decomposition (SVD), principal component analysis (PCA) (Wold et al., 1987) and independent component analysis (ICA) (Comon, 1994).
For SVD, the matrix $\bm{A}$ is decomposed into $\bm{A}=U_{0}\Sigma_{0}V_{0}^{\intercal}$, where $U_{0}\in\mathbb{R}^{n\times n}$ and $V_{0}\in\mathbb{R}^{d\times d}$ are orthonormal matrices, and $\Sigma_{0}\in\mathbb{R}^{n\times d}$ is diagonal. This decomposition reveals the main variability in $\bm{A}$. By retaining the $r\ll d$ largest singular values in $\Sigma$, we approximate $\bm{A}$ as:
@@
where $\bm{U}\in\mathbb{R}^{n\times r}$ contains the leading $r$ columns of $U_{0}$, and $\bm{W}=\Sigma V^{\intercal}\in\mathbb{R}^{r\times d}$ combines the singular values and right singular vectors (see (Eckart and Young, 1936)). SVD is used to capture the most significant patterns in the data by reducing dimensionality while preserving maximum variance.
In PCA and ICA, $\bm{A}$ is projected onto the directions of maximum variance, resulting in a similar approximation:
$\bm{A}\approx\bm{U}\bm{W}\,,$
where $\bm{U}\in\mathbb{R}^{n\times r}$ contains the top $r$ principal components (the eigenvectors of the covariance matrix of $\bm{A}$) for the PCA or contains $r$ independent components for the ICA, and $\bm{W}\in\mathbb{R}^{r\times d}$ contains the corresponding loadings for the PCA or is the mixing matrix for the ICA.
We use PCA to reduce dimensionality by capturing as much variance as possible with the fewest components, making it useful for data compression and noise reduction. And ICA is used to identify underlying factors or sources in the data that are mutually independent, revealing hidden structures not accessible via PCA or SVD.
By expressing $\bm{A}$ as a product of $\bm{U}$ and $\bm{W}$ across these methods, we obtain lower-dimensional representations that capture essential characteristics of the data while emphasizing different statistical properties: SVD focuses on capturing overall variability, PCA emphasizes variance and decorrelation, and ICA highlights statistical independence.
Note that the ICA does not give orthogonal dimensions, as we requested in the decomposition prerequisites, but we decide to study it here anyway, in the hope that the statistical independence between its dimensions may bring interesting results.
For more details on the implementation, we refer the reader to Appendix D.
@@ 4.2 Estimating concept importance with Sobol Indices
We then evaluate the concepts’ importance by applying a feature importance technique on $\bm{G}$ using $c_{g}$, and on $\bm{Y}$ using $c$. More specifically, Sobol indices (Sobol, 1993) are used to estimate the concepts’ importance, as in (Jourdan et al., 2023). A key novelty of the methodology developed in our paper is to simultaneously measure each concept’s importance for the outputs $\bm{G}$ and $\bm{Y}$, and not only $\bm{Y}$ as proposed in (Jourdan et al., 2023).
Figure 2: Co-importance plot for $r=20$ dimensions with respect to Occupation and Gender labels, computed with the Sobol method on RoBERTa model. The color is based on the angle $a=90-\frac{2}{\pi}\arctan{(\frac{y}{x})}$. The angles with extreme values correspond to concepts of high importance for gender but comparatively low importance for occupation. Their removal yields a favorable tradeoff in accuracy/fairness.
To estimate the importance of a concept $\bm{U}_{i}$ for sensitive variable (resp. for the label), we measure the fluctuations of the model’s output $c_{g}(\bm{U}\bm{W})$ (resp. $c(\bm{U}\bm{W})$) in response to perturbations of concept coefficient $\bm{U}_{i}$.
We then propagate this perturbed activation to the model output $\bm{G}=c_{g}(\tilde{\bm{A}})$ (resp. $\bm{Y}=c(\tilde{\bm{A}})$). We can capture the importance that a concept might have as a main effect – along with its interactions with other concepts – on the model’s output by calculating the expected variance that would remain if all the concepts except the $i$-th were to be fixed. This yields the general definition of the total Sobol indices. An important concept will have a large variance in the model’s output, while an unused concept will barely change it.
We explain the total Sobol indices for each concept in detail in Appendix E, along with our perturbation strategy.
@@ 4.3 Neutralizing the embedding with importance-based concept removal
In this final section, we propose to eliminate $k$ concepts that exhibit significant dependence on sensitive variables from the final embedding matrix. This provides us with an improved embedding which can serve as a base to learn a fairer classifier to predict downstream tasks and, thus, a model that is globally more neutral with regard to sensitive variables.
This concept-removal procedure consists of choosing the $k$ concepts that maximize the ratio between the importance for sensitive variable prediction and for the task. This graphically corresponds to searching for the concepts whose angle with respect to the line of equal importance for sensitive variable and task prediction are the highest (see Fig. 2). This strategy allows us to work on a trade-off between the model’s performance and its fairness.
In practice, we sort the concepts $\bm{U}_{1}$,$\bm{U}_{2}$,….$\bm{U}_{r}$ according to this strategy in descending order and delete the first $k$ concepts. This essentially leaves us with a concept base $\tilde{\bm{U}}=(\bm{U}_{k+1},...,\bm{U}_{r})$. This same transformation can be applied to $\bm{W}$’s rows, yielding $\tilde{\bm{W}}$. From the matrix obtained from the matrix product $\tilde{\bm{U}}\tilde{\bm{W}}$, we retrain a classifier $\tilde{c}$ to predict $\bm{Y}$ (refer to Appendix C.2 for more details).
It is interesting to note that, because we are causally removing information from the final latent space, it is impossible for any model to recover it from the truncated space – i.e. we are not limited by the family of models that’s intervening (Ravfogel et al., 2023).
Figure 3: Gender prediction accuracy versus the occupation accuracy drop – using a two-layer MLP – after concept erasure methods on (top left) DistilBERT, (top right) RoBERTa, (bottom left) DeBERTa, and (bottom right) T5 representations. The occupation accuracy drop is reported relatively to the Baseline model, when the value is negative for a method, this means that occupation accuracy is better with the method than it was initially. For INLP and RLACE, points represent different numbers of masked dimensions. For TaCo methods, points represent different numbers of removed concepts. LEACE and the Bios-neutral model are shown as single points, as they do not require parameter adjustments. The horizontal dashed line represents the accuracy of the Optimal Bayes classifier, indicating the theoretical lower bound for gender prediction accuracy. Figures for linear classifier on Figure 7.
@@ 5.1 Dataset and models
To evaluate the efficacy of our method, we employ the Bios dataset (De-Arteaga et al., 2019), which comprises approximately 440,000 biographies annotated with the gender (male or female) of the individuals and their occupations across 28 classes. From this dataset, we derive a second dataset, Bios-neutral, by removing all explicit gender indicators. This Bios-neutral dataset serves as a baseline for comparison. The specifics of the dataset preprocessing are detailed in Appendix B.
For both datasets, we train three encoder transformer models: RoBERTa (Liu et al., 2019a), DistilBERT (Sanh et al., 2019), and DeBERTa (He et al., 2020), and an encoder-decoder transformer model: T5 (Raffel et al., 2020), to perform the occupation prediction task. The training procedures and hyperparameter settings for these models are provided in Appendix C.1.
5.2 Quantitative Evaluation of Concept Erasure
We compare our method against several state-of-the-art concept erasure techniques:
$\bullet$ Iterative Nullspace Projection (INLP) (Ravfogel et al., 2020)
$\bullet$ Reduced Linear Adversarial Concept Erasure (RLACE) (Ravfogel et al., 2022)
$\bullet$ Linear Adversarial Concept Erasure (LEACE) (Belrose et al., 2023).
Additionally, we include:
$\bullet$ Baseline Model: The original model without any modifications.
$\bullet$ Bios-neutral Model: A model similar to the baseline but trained on the Bios-neutral dataset.
$\bullet$ Optimal Bayes: corresponding to the optimal Bayes classifier $\mathbb{P}(G|Y)$ predicting gender from occupation.
$\bullet$ TaCo Variants: Our proposed method using different decompositions—Singular Value Decomposition (SVD), Principal Component Analysis (PCA), and Independent Component Analysis (ICA).
Figures 3 and 7 provide a comparative analysis of the debiasing methods, with the x-axis showing the drop in occupation prediction accuracy relative to the baseline model, and the y-axis representing the gender prediction accuracy. In these figures, the x-axis is truncated at a 0.1 accuracy drop, as losing more than 10% in task accuracy is not particularly informative. The full versions of these plots, without the x-axis truncation, are provided in Figures 8 and 9, on appendix.
The varying points for each method correspond to different parameter settings:
$\bullet$ INLP and RLACE: We manipulate the number of dimensions masked from the final latent representation (up to 768 dimensions for BERT-type representations and 512 dimensions for T5 representation). Specifically, for INLP, we masked [200, 300, 400, 500, 550, 600, 650, 700, 750] dimensions for BERT-type representations and [100, 200, 300, 350, 400, 450, 500] dimensions for T5 representation, and for RLACE, we masked [1, 5, 10, 100, 200, 500] dimensions for BERT-type representations and [1, 5, 10, 100, 200, 300] dimensions for T5 representation. RLACE typically requires fewer dimensions to be hidden to achieve effectiveness (Ravfogel et al., 2022).
$\bullet$ TaCo Variants: For each decomposition (SVD, PCA, ICA), we vary the number of concepts (i.e., dimensions in the decomposition’s basis) removed, ranging from 1 to $r-1$, where $r=20$ is the number of components used. This results in 19 data points for each TaCo variant.
$\bullet$ LEACE: Since LEACE does not require parameter tuning, it is represented by a single point.
$\bullet$ Baseline and Bios-neutral Models: Each is depicted by a single point corresponding to their respective accuracies.
$\bullet$ Optimal Bayes: Represented by a horizontal dashed line on the graph, as it has no associated occupation accuracy value. This line indicates the theoretical limit for gender prediction accuracy, below which further reduction is not meaningful.
All results are averaged over five independent runs to ensure robustness.
We assess the $\nu$-information in both linear and non-linear settings: linear $\nu$-information – measured by the accuracy of a logistic regression classifier predicting the sensitive attribute and the label –, and non-linear $\nu$-information – measured using a two-layer Multilayer Perceptron (MLP) classifier. Detailed results of these classifiers are provided in Appendix C.2.
As illustrated in Figure 7 (in Appendix), LEACE and RLACE excel at obscuring sensitive information from linear classifiers, aligning with their design objectives. Notably, TaCo with PCA closely matches INLP’s performance in the linear context, despite INLP’s optimization for linear scenarios.
In contrast, Figure 3 highlights TaCo’s superiority in non-linear settings. While LEACE and RLACE effectively conceal gender information from linear models, they fail to prevent non-linear models like MLPs from retrieving this information. TaCo, particularly with PCA and SVD decomposition, significantly reduces the gender information accessible to non-linear classifiers while maintaining high occupation prediction accuracy. This demonstrates a more favorable trade-off compared to simply removing explicit gender indicators, as done in the Bios-neutral model.
Further analysis reveals that TaCo with PCA slightly outperforms TaCo with SVD in non-linear contexts, especially in scenarios where task accuracy remains largely unaffected. This suggests that PCA is the optimal decomposition for TaCo when aiming to balance task performance with sensitive information removal. Conversely, TaCo with ICA performs less effectively across both linear and non-linear settings.
In summary, our proposed TaCo method, especially when combined with PCA decomposition, effectively mitigates the leakage of sensitive information in both linear and non-linear models. It achieves this with minimal impact on task performance, outperforming existing state-of-the-art techniques and offering a robust solution for bias mitigation in language models.
5.3 Interpretability of Concepts Through Explainability Technique
Beyond effectively erasing sensitive information from embeddings against non-linear adversaries, our method offers an additional advantage in terms of interpretability. Due to its structure, which parallels practices in explainable AI (XAI), TaCo facilitates the straightforward application of interpretability techniques to the identified concepts. Specifically, we applied the COCKATIEL method (Jourdan et al., 2023) to interpret the concepts derived from our matrix decompositions before their removal. This analysis, detailed in Appendix G, demonstrates that concepts obtained via SVD are semantically meaningful and provide insights into how the model processes information related to both the sensitive attribute and the primary task. However, the interpretability of these concepts varies according to the decomposition method used, so this must be taken into account when selecting the decomposition.
This interpretability aspect underscores the transparency of our method and highlights its potential for enhancing model accountability, an advantage that may not be as readily achievable with other concept erasure techniques.
@@ 6 Conclusion
In this work, we introduced TaCo, a novel concept erasure method that removes sensitive information from latent space representations, ensuring fairness against both linear and non-linear adversaries. Our method involves concept discovery, ranking based on importance to both the sensitive attribute and target label, and removing concepts that affect the sensitive variable while preserving task performance.
Experiments on the Bios dataset showed that TaCo outperforms state-of-the-art methods, effectively reducing non-linear classifiers’ ability to predict sensitive attributes without compromising model accuracy.
Additionally, TaCo enhances interpretability by enabling insights into the factors influencing sensitive attribute predictions. By applying explainability techniques like COCKATIEL, we contribute to both model transparency and accountability.
Our findings highlight the importance of addressing non-linear dependencies in mitigating biases, positioning TaCo as a significant step toward fairer and more transparent AI systems.
@@
While our proposed method demonstrates effectiveness in mitigating the influence of sensitive attributes, it does not guarantee a clear separation between concepts containing information about the output variable and those related to the sensitive attribute. In some cases, the concepts identified through decomposition may be inherently entangled, simultaneously encoding information pertinent to both the task and the sensitive variable. This entanglement poses a challenge in precisely isolating and removing only the unwanted information without adversely affecting the model’s performance on the primary task.
Moreover, our approach to interpreting concepts post-removal has room for improvement. Interpreting concepts in natural language processing remains a complex endeavor due to the intricate and high-dimensional nature of textual data. Developing more sophisticated methods for concept interpretation is essential and warrants dedicated future research.
Another limitation pertains to the scope of our application. While we focused on mitigating gender bias using the Bios dataset, which considers gender as a binary attribute, we acknowledge that gender is inherently non-binary. Consequently, models trained on such datasets may not fully capture the spectrum of gender identities present in real-world scenarios. Although our method is designed to operate effectively with variables encompassing multiple classes, including non-binary gender categorizations, the lack of datasets reflecting this diversity limits the evaluation of our approach in more inclusive settings. This underscores the need for the NLP community to develop and adopt datasets that incorporate non-binary and diverse representations of sensitive attributes, thereby promoting fairness and inclusivity in AI applications.
Ethics Statement
With the advent of LLMs and their expanding applications, it is crucial to ensure that biases are adequately addressed before deploying these models in high-stakes applications. In this paper, we introduce a method that not only mitigates biases but does so in a manner that promotes human understanding of the specific biases removed from the model. Transparency in identifying bias sources and understanding the trade-off involved are vital aspects of our approach.
An important consideration is the computational efficiency of our method, which significantly reduces the resource burden. Our approach only requires training a classification head while keeping the backbone of the model unchanged. This computational affordability makes our method particularly valuable for companies with limited budgets, enabling them to effectively address bias concerns without incurring excessive costs.
We believe that it is our ethical responsibility to prioritize fairness and accountability in the development and deployment of AI systems. By providing a debiasing technique that offers interpretability, computational efficiency, cost-effectiveness and ease of use, we aim to support the responsible use of LLMs, ensuring they are better aligned with ethical standards and advancing the progress of unbiased AI technologies.